In how many ways can 12 children be placed on 3 different teams of 3, 5, and 4 members?
Answer by AnlytcPhil(1806)

C(12,3)×C(9,5)×C(4,4) = 220×126×1 = 27720

Notice that it won't matter what order you place them on the teams,
because you get the same answer, regardless:

C(12,3)×C(9,4)×C(5,5) = 220×126×1 = 27720

C(12,5)×C(7,3)×C(4,4) = 792×35×1 = 27720

C(12,5)×C(7,4)×C(3,3) = 792×35×1 = 27720

C(12,4)×C(8,3)×C(5,5) = 495×56×1 = 27720

C(12,4)×C(8,5)×C(3,3) = 495×56×1 = 27720

Edwin